OpenStudy (anonymous):

if the product of four cubed and y is 320 then y =

OpenStudy (bahrom7893):

4^3*y=320 64y=320 y = 320/64 = 5

OpenStudy (bahrom7893):

Please click on Good Answer if I helped, thanks!